OpenStudy (anonymous):

Help with a simplied expression y=20 triangle height = 2/5y+4 base=1/2y

OpenStudy (anonymous):

y = 20 height = 2/5 *(20+4) = 2/5 * 24 = 48/5 = 9.6 Base = 1/2*20 = 10

OpenStudy (anonymous):

the teacher wants an expression?

OpenStudy (anonymous):

expression for area ?

OpenStudy (anonymous):

a simplified expression for area is all it says

OpenStudy (anonymous):

Area = 1/2 base times height

OpenStudy (anonymous):

so half of 10 times 12

OpenStudy (anonymous):

area = 1/2 * base * height = 1/2 * 2/5 * (y+4) * 1/2* y area = y*(y+4)/ 10 area = 1/10 * (y^2 +4y) Plug in y = 20 area = 1/10 * (400 + 80) = 1/10 * 480 = 48

OpenStudy (anonymous):

which is 60

OpenStudy (anonymous):

ill try it

OpenStudy (anonymous):

wait - is it 2/5y or (2/5)y?

OpenStudy (anonymous):

its 2/5y

OpenStudy (anonymous):

ok so forget my answer

OpenStudy (anonymous):

so what do itype as my answer?

OpenStudy (anonymous):

Try Rita's

OpenStudy (anonymous):

0k